### v1

# SAFE: Multitask Failure Detection for Vision-Language-Action Models

Abstract
While vision-language-action models (VLAs) have shown promising robotic behaviors across a diverse set of manipulation tasks, they achieve limited success rates when deployed on novel tasks out-of-the-box. To allow these policies to safely interact with their environments, we need a failure detector that gives a timely alert such that the robot can stop, backtrack, or ask for help. However, existing failure detectors are trained and tested only on one or a few specific tasks, while VLAs require the detector to generalize and detect failures also in unseen tasks and novel environments. In this paper, we introduce the multitask failure detection problem and propose SAFE, a failure detector for generalist robot policies such as VLAs. We analyze the VLA feature space and find that VLAs have sufficient high-level knowledge about task success and failure, which is generic across different tasks. Based on this insight, we design SAFE to learn from VLA internal features and predict a single scalar indicating the likelihood of task failure. SAFE is trained on both successful and failed rollouts, and is evaluated on unseen tasks. SAFE is compatible with different policy architectures. We test it on OpenVLA, $\pi_{0}$, and $\pi_{0}$-FAST in both simulated and real-world environments extensively. We compare SAFE with diverse baselines and show that SAFE achieves state-of-the-art failure detection performance and the best trade-off between accuracy and detection time using conformal prediction. More qualitative results can be found at https://vla-safe.github.io/.
1 Introduction
Recently, scaling up robot manipulation datasets has enabled the development of large vision-language-action (VLA) models, which are generalist manipulation policies that can follow language instructions and accomplish a wide range of tasks [1, 2, 3, 4, 5, 6]. However, when VLAs are directly deployed on unseen tasks without collecting additional demonstrations and finetuning the model, they still suffer from limited success rates and a wide range of failure modes. This has been demonstrated by evaluations in recent work [2, 7, 4]: while their success rates on seen tasks are as high as 80-90%, those on unseen ones drop to 30-60% out-of-the-box. Therefore, to safely and reliably deploy VLA policies in the real world, it is important to promptly detect their potential failures.
Most existing failure detection methods train a separate failure detector for each task, and evaluate the detector only on that task [8, 9, 10, 11, 12, 13, 14, 15, 16]. While these methods work well for specialist policies, they do not suit generalists like VLAs. VLAs are designed to accomplish diverse tasks and may frequently encounter novel task instructions and unseen environments during deployment. In such cases, it is impractical to exhaustively collect rollouts and train a failure detector for every new task. Some recent works introduce task-generic failure detectors, but they either require sampling multiple actions [17] or need to query a large VLM [18, 19], which poses significant inference overhead for large VLAs in the real world. This motivates the need for an efficient and multitask failure detector that can generalize to unseen tasks zero-shot and detect failures in a timely manner during the on-policy rollout of the VLA.
In this paper, we focus on the multitask failure detection problem. This setting evaluates failure detection performance of a VLA policy without collecting rollouts or finetuning the failure detector on unseen tasks. To our knowledge, such multitask failure detection ability for VLAs has not been shown in the literature. To tackle this problem, we study the internal features of VLAs and find that they capture high-level knowledge about task success and failure. As shown in Fig. 1, such knowledge can effectively distinguish failed rollouts from successful ones and is generic across different tasks.
Figure 1: The VLA’s internal features capture high-level information about task success and failure. When the VLA is failing, even though from different tasks, the features fall in the same “failure zone”. This motivates SAFE, an efficient multitask failure detector that is based on VLA internal features and can generalize to unseen tasks. Plot (a) visualizes latent features of $\pi_{0}$-FAST on LIBERO-10 [20] using t-SNE [21]. For successful rollouts, their features are colored in blue; for failed rollouts, features are colored in the blue-red gradient based on timesteps. Plot (b) visualizes the same set of t-SNE features, colored by task ID. In (c), we show two example rollouts over time and mark their corresponding projected features in (a) and (b).
Based on this insight, we introduce SAFE, a ScAlable Failure Estimation method that scales across diverse tasks for generalist policies like VLAs. SAFE takes in a VLA’s internal features and regresses them to a single scalar indicating the likelihood of failure. By training on successful and failed rollouts of multiple tasks, SAFE learns to identify task-generic representations for failure detection. To determine the threshold for failure detection, we adopt the functional conformal prediction (CP) [22, 8] framework and calibrate the prediction band on the seen tasks. We conduct failure detection experiments on OpenVLA [2], $\pi_{0}$ [4] and $\pi_{0}$-FAST [5], in both simulation and the real world. For evaluation, we adapt diverse baseline failure detection methods from both the LLM literature [23, 24] and the robot learning literature [8, 17] onto VLAs. SAFE and baselines are evaluated on both training tasks and a set of held-out tasks. Experiments show that SAFE outperforms other existing baselines and achieves the best trade-off between accuracy and timeliness for failure detection. The contributions of our paper can be summarized as follows:
  • We analyze the VLA feature space and show that across different task instructions and environments, VLA’s internal features are separated for successful and failed rollouts.
  • We propose SAFE, a multitask failure detector designed for generalist robot policies. By operating on latent features, training on multiple tasks, and using conformal prediction methods, SAFE shows generalization ability in detecting failures on unseen tasks.
  • We evaluate SAFE and diverse baselines on several recent large VLA models in both simulation and the real world. Experiments show that SAFE outperforms baselines and achieves state-of-the-art (SOTA) performance.
2 Related Work
2.1 Vision-Language-Action Models
Recent advances in large-scale machine learning and the availability of extensive robot demonstration datasets have paved the way for VLA models [1, 2, 3, 4, 7, 25, 26, 6]. These generalist robotic policies are initialized from pretrained large-scale VLMs [27, 28, 29], and thus inherit the strong ability to understand diverse semantic concepts from both images and language. They are augmented with an action head that produces continuous control signals, through per-step binning [1, 7, 2, 25], diffusion networks [3, 4, 30, 31] or frequency-space tokenization [5]. These VLAs are then trained on vast robotic datasets covering a wide array of tasks [32, 33, 34]. As a result, VLAs can successfully perform familiar tasks in new environments and even tackle previously unseen tasks when provided with novel language instructions. Nevertheless, significant variability in real-world deployments and the challenging domain gaps between training and testing environments continue to hinder VLA performance. Most state-of-the-art VLA models achieve success rates between 30% and 60% when evaluated out-of-the-box on real robots with unseen task instructions [2, 32, 4]. These limitations highlight the need for robust multitask failure detection methods tailored to generalist VLA models.
2.2 Failure Detection in Robot Manipulation
Monitoring failures is critical when deploying robotic policies in real-world environments, as even minor errors can result in hazardous conditions [35, 36, 37]. The literature on failure detection in robot learning can be broadly divided into unsupervised out-of-distribution (OOD) detection [8, 9, 10, 11] and supervised failure detection [9, 12, 13, 14, 15, 16]. OOD detection-based methods treat successful executions as the in-domain baseline and consider any deviation from this norm as a failure. However, the assumption that any unseen scenario constitutes a failure is overly restrictive for generalist VLAs, which may frequently encounter unseen tasks at test time. These unseen tasks are likely different from the in-domain training data but should not be simply treated as failures. Our proposed method, SAFE, falls within the supervised failure detection category, leveraging both successful and failed rollouts to train a failure classifier. Yet SAFE trains a single unified failure detector for all tasks, while other existing ones typically require training and calibration of separate classifiers for each task and have not yet demonstrated effectiveness on generalist policies like VLAs. Some recent works have explored the multitask failure detection task by designing action consistency scores [17] or instruction-finetuning a VLM [18, 19], but they require either sampling multiple actions or querying a large VLM, which poses significant overhead for controlling robots in real time.
Recently, FAIL-Detect [8] conducted a systematic evaluation of various failure detection methods, including OOD detection-based approaches [38, 39, 9], smoothness-based techniques [40], and consistency-based strategies [17]. Their experiments indicate that the best performance was achieved by LogpZO [38] and RND [39], both using a learned network to model data distribution and detect OOD states. However, their evaluation is limited to only single-task policies, and our evaluation in multitask setting shows that their best performing RND and LogpZO methods suffer from overfitting to the training tasks.
2.3 Uncertainty Quantification for LLM
Although LLMs and VLMs have demonstrated remarkable understanding and generative capabilities across various tasks, they are prone to producing hallucinated responses [41, 42, 43]. Numerous methods have been developed for uncertainty quantification (UQ) in LLMs/VLMs. Token-level uncertainty quantification methods estimate uncertainty by analyzing the probability distribution over each generated token to assess the likelihood of an entire response [44, 45, 46]. In contrast, semantic-similarity methods generate multiple responses to the same query and evaluate their semantic alignment [24, 47, 48]; a higher variance among responses typically signals low confidence. Since vision-language-action models (VLAs) share the generative nature and transformer architecture of LLMs/VLMs, we adapt these UQ methods to VLAs as promising baselines and evaluate their performance on failure detection. Recent research has also explored the internal latent space of LLMs for hallucination detection [49, 50, 51, 52, 53, 54, 55]. These methods train a classifier on internal latent features to distinguish between truthful and hallucinated outputs, paralleling supervised failure detection techniques in robotics. This approach has proven to be simple, efficient, and effective for UQ in LLMs. In our study, we investigate its application to large VLA policies and observe promising performance in robotic tasks.
3 Problem Formulation
This work aims to detect when a robot policy fails its task execution. Specifically, we aim for a multi-task failure detector that performs well when generalist VLAs perform novel tasks at inference time. At timestep $t$, a VLA is given an input observation $\mathbf{o}_{t}$, consisting of RGB images, natural language instruction, and current robot state, and outputs a control signal $\mathbf{A}_{t}=[\mathbf{a}_{t},\mathbf{a}_{t+1},\dots,\mathbf{a}_{t+H-1}]$, which is a chunk of actions for the next $H$ timesteps. The first $H^{\prime}$ ($H^{\prime}\leq H$) actions in $\mathbf{A}_{t}$ are executed, and then the VLA replans a new action sequence $\mathbf{A}_{t+H^{\prime}}$ at time $t+H^{\prime}$. We denote the internal embedding vector within the VLA model at time $t$ as $\mathbf{e}_{t}$. Some VLAs [1, 2, 4, 5] also decode a series of $m$ tokens $\mathbf{W}_{t}=[\mathbf{w}_{t}^{1},\dots,\mathbf{w}_{t}^{m}]$ before converting them into the actual action vector. To train and evaluate failure detection models, we run the VLA on different tasks in simulation or the real world, collect the rollout trajectory $\tau_{i}=\{(\mathbf{o}_{t},\mathbf{e}_{t},\mathbf{W}_{t},\mathbf{A}_{t})\}_{t=0,H^{\prime},\dots,nH^{\prime}}$ with time duration $T=nH^{\prime}$, and annotate each rollout with a failure label $y_{i}$ ($y_{i}=1$ if the robot fails to accomplish the task and $y_{i}=0$ if the robot succeeds). Note that for training, we only use the trajectory-level annotation $y_{i}$, and don’t require knowing the exact timestep when the policy starts to fail. A failure detector receives the rollout information up to time $t$ and predicts a failure score $s_{t}$, indicating the likelihood of task execution failure at time $t$. If $s_{t}$ exceeds a threshold $\delta_{t}$, a failure flag is raised, and then either the task execution is aborted or a human monitor will step in and take over the control. In this work, we use conformal prediction [56] to calibrate the threshold $\delta_{t}$.
In experiments, we split all tasks into seen and unseen subsets, where rollouts from seen tasks are used for training $\mathcal{D}_{\text{train}}$ and validation $\mathcal{D}_{\text{eval-seen}}$, and all rollouts from unseen tasks $\mathcal{D}_{\text{eval-unseen}}$ are reserved for testing the cross-task generalization ability of failure detectors. Failure detectors are trained on $\mathcal{D}_{\text{train}}$, and evaluated on $\mathcal{D}_{\text{eval-seen}}$ for hyperparameter tuning and in-domain performance, and tested on $\mathcal{D}_{\text{eval-unseen}}$ for out-of-distribution generalization.
4 Method
4.1 Visual Analysis on VLA Latent Space
VLAs process multi-modal inputs and extract rich semantic information in their internal feature space. We hypothesize that these features also capture the high-level and abstract knowledge about task execution success/failure, by separating features from successful/failed rollouts into different regions. We study this hypothesis by visualizing the VLA features in Fig. 1, where we plot the internal features from $\pi_{0}$-FAST[5] when running the LIBERO-10 benchmark [20]. Fig. 1(a) demonstrates that when the VLA is failing, its internal features are grouped in the same region in the feature space (“failure zone”). Comparing Fig. 1(a) and Fig. 1(b), we can further see that although the features are extracted from different tasks with various instructions, objects and environments, when the VLA fails, its features fall in the same “failure zone”. Fig. 1(c) further illustrates how VLA’s features evolve in the feature space when VLA progresses temporally. From Fig. 1(c), we can see that failure rollout initially stays out of the “failure zone” when it progresses normally, and when the robot mistakenly drops the pot in the middle of execution and starts to fail, it steps into the “failure zone”. On the contrary, for the successful rollout, its features always stay out of the “failure zone”.
This visual analysis shows that the VLA’s internal features for succeeding and failing task executions are well separated in the feature space, and this separation is general across different tasks. Furthermore, during task execution, the features reflect how well the VLA performs on the current tasks in a timely manner. Inspired by this observation, we design SAFE, which uses the internal features of VLAs for failure detection.
Table 1: Failure detection results on simulation benchmarks, measured by area under ROC (ROC-AUC). “-” indicates that the failure detection method does not apply. Entries with gray background indicate the failure detection methods that sample 10 actions per inference timestep, while others use only 1 action. The first and second best-performing methods are colored in red and orange, respectively. Results are averaged over 3 random seeds with different splits of seen and unseen tasks.
[…]
4.2 Failure Detection by Feature Probing
We design SAFE to learn the abstract information from the VLA’s internal features and determine whether the task execution is failing. We extract the VLA’s hidden state vectors from the final layer, before being decoded to token logits [2, 5] or a velocity field [4]. We ablate different ways to aggregate the internal features into a single embedding vector $\mathbf{e}$, and select the best one based on $\mathcal{D}_{\text{eval-seen}}$ performance. Please refer to Appendix for details on VLA feature extraction.
The failure detector $f(\mathbf{e}_{0:t})$ takes as input the VLA’s features $\mathbf{e}_{0:t}=\{\mathbf{e}_{1},\dots,\mathbf{e}_{t}\}$ up to the current timestep $t$, and is trained to predict $s_{t}$. We explore the two backbone designs for SAFE: a multi-layer perceptron ($f_{\text{MLP}}$) and an LSTM [57] ($f_{\text{LSTM}}$). Both models are designed to be simple (only one or two layers), to avoid overfitting and improve generalization ability on unseen tasks. For $f_{\text{MLP}}$, we use an MLP $g(\cdot)$ to project $\mathbf{e}_{t}$ into a single scalar for each timestep $t$ independently and accumulate the outputs as the failure score, i.e. $f_{\text{MLP}}(\mathbf{e}_{0:t})=\sum_{\tau=1,\dots,t}\sigma(g(\mathbf{e}_{\tau})),$ where $\sigma(\cdot)$ is a sigmoid function and therefore $0<s_{t}<t$. To train the MLP model, we apply an L1 loss on all timesteps to push up the scores for failed rollouts and push down those for successful ones. Specifically, $L_{\text{MLP}}=\sum_{i}\left[y_{i}\sum_{t}(t-s_{t})+(1-y_{i})\sum_{t}s_{t}\right],$ where index $i$ iterates over all data points in $\mathcal{D}_{\text{train}}$.
For $f_{\text{LSTM}}$, we use an LSTM model to sequentially process the input stream of VLA’s features $\mathbf{e}_{0:t}$ and project the hidden state vector of LSTM into a scalar score. Specifically, $f_{\text{LSTM}}(\mathbf{e}_{0:t})=\sigma(\text{LSTM}(\mathbf{e}_{0:t})),$ where a sigmoid function $\sigma(\cdot)$ is applied to normalize the output score s.t. $0\leq s_{t}\leq 1$. To train the LSTM model, we apply a binary cross entropy loss on all timesteps, i.e. $L_{\text{LSTM}}=\sum_{i}\sum_{t}\left[y_{i}\log(s_{t})+(1-y_{i})\log(1-s_{t})\right].$
4.3 Threshold Selection by Conformal Prediction
When the predicted failure score $s_{t}$ exceeds the time-varying threshold $\delta_{t}$, we raise a failure flag. To determine $\delta_{t}$ in a principled way, we adopt the functional conformal prediction (CP) framework [22]. Functional CP constructs a time-varying prediction band by leveraging the distribution of $s_{t}$ observed in successful rollouts within a calibration set. Under the exchangeability assumption [58] and given a user-specified significance level $\alpha$, the CP band guarantees that, for a new successful rollout, its $s_{t}$ will lie within this band at all times $t$ with probability $1-\alpha$. Conversely, if the score of a test rollout exits the band at time $t$, we can declare a failure with nominal confidence $1-\alpha$.
Formally, given a time series of any scalar score $s_{t}$ and a user-specified significance level $\alpha\in(0,1)$, functional CP gives a distribution-free prediction band $C_{\alpha}$. Following [8], we adopt the one-sided time-varying CP band formulation, where $C_{\alpha}$ is a set of intervals $\{[\text{lower}_{t},\text{upper}_{t}]:t=1,\dots,T\}$, where $\text{lower}_{t}=-\infty$ and $\text{upper}_{t}=\mu_{t}+h_{t}$, with a time-varying mean $\mu_{t}$ and a bandwidth $h_{t}$. This band is calibrated on successful rollouts in $\mathcal{D}_{\text{eval-seen}}$. Under mild assumptions [59, 60], for any new successful rollout, $s_{t}<\mu_{t}+h_{t}$ holds for all $t=1,\dots,T$ with probability $1-\alpha$. Intuitively, this gives a guarantee that the false positive rate of the failure detector (a failure flag is raised at any time during a successful rollout) is at most $\alpha$. We use $\text{upper}_{t}$ as the failure flag threshold $\delta_{t}$, and more details about functional CP can be found in Appendix.
5 Experiments
Figure 2: Illustration of real-world experiment setup (left) and example rollouts collected (right).
5.1 Evaluation Benchmarks
LIBERO [20]: The LIBERO benchmark has been widely adopted for evaluating VLA models in simulation [2, 4, 5, 6]. Among the LIBERO task suites, the LIBERO-10 suite consists of 10 long-horizon tasks with diverse objects, layouts, and instructions, and is considered the most challenging one. Therefore, we use LIBERO-10 in our experiments and test OpenVLA [2], $\pi_{0}$ [4] and $\pi_{0}$-FAST [5] on it. We adopt the model checkpoints that are finetuned on the LIBERO benchmark and released by their authors. In experiments, 3 out of 10 tasks are randomly picked and reserved as unseen tasks.
SimplerEnv [61]: SimplerEnv provides a high-fidelity simulation environment for manipulation policies, which are replicas of the demonstration data from RT-series [1, 7, 32] and BridgeData V2 [34]. On SimplerEnv, we test pretrained $\pi_{0}$ models from a reproduction [62], which we denote as $\pi_{0}^{*}$ in this paper. We train and evaluate the failure detection methods on the Google Robot embodiment [1] and on the WidowX embodiment [34], respectively. We exclude the “ pick up coke” task because $\pi_{0}^{*}$ rarely fails on it (success rate at 98%). This leaves 4 tasks for each embodiment, among which 3 tasks are seen and 1 task is unseen.
Real-robot Experiments: We deploy the $\pi_{0}$-FAST-DROID checkpoint [^1] https://github.com/Physical-Intelligence/openpi provided by [4, 5] on a Franka Emika Panda Robot. This checkpoint has been finetuned on the DROID dataset [33], and we do not further collect demonstrations or finetune the VLA model. We design 13 tasks and collect 30 successful and 30 failed rollouts for each task. The real-robot setup and example rollouts are visualized in Fig. 2. In experiments, 3 tasks out of 13 are randomly selected as unseen tasks.
Figure 3: Performance of functional CP by varying significant level $\alpha$ on $\mathcal{D}_{\text{eval-unseen}}$. The plots show the change of balanced accuracy (bal-acc) w.r.t. average detection time (T-det). Good failure detection methods should detect policy failures both accurately (high bal-acc) and proactively (lower T-det), and thus place curves towards the top left in each plot. We can see in all simulation experiments, the proposed SAFE-LSTM and SAFE-MLP perform better or on par with the best baselines. Note that baselines in gray require multiple action samples.
5.2 Uncertainty Quantification Baselines
Estimating uncertainty in generated responses has been widely used to detect truthfulness or hallucination in LLMs [23, 24, 63, 43]. For VLAs, uncertainty in the generated actions may indicate a lack of ability to solve the given task, and thus correlates with task failures. Therefore, we first adapt the UQ methods from the LLM literature to VLAs and use them as failure detection baselines.
Token uncertainty-based methods aggregate the predictive uncertainty from each generated token. These methods are efficient, as they only require a single forward inference. Given the generated tokens $\mathbf{W}_{t}=[\mathbf{w}_{t}^{1},\dots,\mathbf{w}_{t}^{m}]$, we denote the probability of sampling the token $\mathbf{w}_{t}^{i}$ as $p_{i}$ and the entropy over the distribution of the $i^{\text{th}}$ token as $H_{i}$. We adopt the token-based uncertainty estimation methods proposed in [23] as follows:
Token max prob.: $\max_{i}(-\log p_{i})$; Token avg prob.: $-\frac{1}{m}\sum_{i}\log p_{i}$;
Token max entropy: $\max_{i}H_{i}$; Token avg entropy: $\frac{1}{m}\sum_{i}H_{i}$.
Sample consistency-based methods estimate uncertainty as the inconsistency within multiple generated sentences [23, 24, 63]. For VLA models, the output actions are continuous vectors, and we can measure inconsistency by their variance. Specifically, at time $t$, given $K$ sampled actions $\mathcal{A}_{t}=\{\mathbf{A}_{t}^{k}\}_{k=1,\dots,K}$, we measure the uncertainty as the total variation over the set of vectors: action total var. = trace(cov($\mathcal{A}_{t}$)). Similarly, we also compute variation for the translational (trans. total var.), rotational (rot. total var.), and gripper control (gripper total var.) components of $\mathcal{A}_{t}$.
[…]
5.3 Failure Detection Baselines
Embedding Distance: We compare to baselines that directly use the distances in the feature space as failure scores. Specifically, instead of training a neural network, all VLA embeddings from $\mathcal{D}_{\text{train}}$ are stored in the two feature sets, $E_{\text{succ}}$ and $E_{\text{fail}}$, containing all VLA embeddings from successful and failed rollouts respectively. During evaluation on $\mathcal{D}_{\text{eval-seen}}$ and $\mathcal{D}_{\text{eval-unseen}}$, failure scores are computed as $s_{t}=d(\mathbf{e}_{t},E_{\text{succ}})-d(\mathbf{e}_{t},E_{\text{fail}})$, where $d(\cdot,\cdot)$ measures the distance between a single vector and a set of vectors. Intuitively, if $\mathbf{e}_{t}$ is far from $E_{\text{succ}}$ and close to $E_{\text{fail}}$, it’s more likely to fail. Following [17, 10], we ablate different types of distance, including Mahalanobis distance, and Euclidean and Cosine distance averaged over $k$-Nearest Neighbors of $\mathbf{e}_{t}$. We also compare to the PCA-KMeans distance measure from [9].
Learned OOD Detector: We adopt RND [39] and LogpZO [8], which are shown to be the best-performing OOD detection-based failure detectors by [8]. Both methods use a neural network $f^{\text{OOD}}(\cdot)$ to model the embedding distribution from successful rollouts and return an OOD score for a new embedding. We adapt them to learn from both successful and failed rollouts by training two models, $f^{\text{OOD}}_{\text{succ}}(\cdot)$ and $f^{\text{OOD}}_{\text{fail}}(\cdot)$, on $E_{\text{succ}}$ and $E_{\text{fail}}$ respectively. Similar to embedding distance baselines, the failure score is computed as $s_{t}=f^{\text{OOD}}_{\text{succ}}(\mathbf{e}_{t})-f^{\text{OOD}}_{\text{fail}}(\mathbf{e}_{t})$.
[…]
5.4 Evaluation Protocol
We consider two types of evaluation. The first type evaluates how well $s_{t}$ separates the successful and failed rollouts across all possible selections of $\delta_{t}$. Following the evaluation protocol widely adopted in the LLM UQ literature [49, 53, 23, 65], we use the area under ROC curve (ROC-AUC) metric. Furthermore, because a failure flag is raised whenever $s_{t}$ exceeds $\delta_{t}$, a successful rollout (ground truth negative) becomes a false positive whenever $s_{t}>\delta_{t}$, and remains a true negative only if $s_{t}\leq\delta_{t}$ for all time. Therefore, we consider the max-so-far score $\bar{s}_{t}=\max_{\tau=1,\dots,t}s_{\tau}$ and compute the ROC-AUC metric based on $\bar{s}_{T}$, the maximum failure score throughout the entire rollout.
[…]
6 Results
Figure 4: Failure scores predicted by SAFE-LSTM and the corresponding camera observations. The green areas show the functional CP band $C_{\alpha}$. Once failure scores exceed $C_{\alpha}$, a failure flag is raised. In (a), $\pi_{0}$-FAST policy misses the insertion, and its actions become unstable after that. In (b) and (c), OpenVLA and $\pi_{0}^{*}$ miss the grasp but still proceed to the placing action, causing a failure detection. Note that these tasks are not seen when training SAFE-LSTM.
6.1 How well do failure detectors distinguish failures from successes?
In Table 1 and Fig. 5 (a), we report the ROC-AUC metric based on $\bar{s}_{T}$, in simulation and real-world experiments, respectively. With a higher ROC-AUC metric, a failure detector achieves higher accuracy averaged over all possible thresholds. The tables show that Token Unc. methods have poor performance, which is aligned with findings in the LLM literature [24, 23]. On the other hand, the Sample Consist. and STAC [17] methods, which require multiple action samples, perform better and even achieve the best performance on unseen tasks in $\pi_{0}$-FAST LIBERO (STAC) and $\pi_{0}$ LIBERO (Gripper total var.). However, as these methods require multiple action samples, they cause significant overhead for VLA models and thus are not applicable to real robots. Embed. Distr. methods perform well, achieving the best performance in two simulation benchmarks ($\pi_{0}$ and $\pi_{0}$-FAST) and are the second best in the real world. This demonstrates that a VLA’s internal features are informative about task execution success/failure. The proposed SAFE methods perform better or on par with the best baselines, consistently in all settings. Averaged across simulation benchmarks, SAFE-MLP and SAFE-LSTM have similar performance, both outperforming the best baseline by 4-5% on unseen tasks, while still achieving the best performance on seen tasks. On the real robot, SAFE-MLP achieves the best performance and SAFE-LSTM performs closely with the best baseline (Euclid. $k$-NN). Comparing SAFE with Embed. Distr. methods, we attribute the success of SAFE to its stronger ability to extract high-level abstract information from raw feature vectors through learned neural networks.
Task Split Seen Unseen Max prob. 53.74 48.59 Avg prob. 51.60 47.30 Max entropy 59.23 53.50 Avg entropy 50.67 46.08 Mahala. dist. 75.54 53.93 Euclid. $k$-NN 80.35 60.27 Cosine. $k$-NN 80.23 59.51 PCA-KMeans 49.98 51.03 RND 62.00 45.83 LogpZO 64.43 52.24 STAC-Single 45.24 38.01 SAFE-LSTM 77.27 58.70 SAFE-MLP 86.76 64.16 (a) Failure Detection ROC-AUC
Figure 5: Quantitative results (a) and qualitative examples (b-e) from SAFE-MLP in the real world experiments. The ROC-AUC results are averaged over 5 random seeds with different task splits.
6.2 How do detection accuracy and detection time trade off using functional CP?
In Fig. 3, we use $\mathcal{D}_{\text{eval-seen}}$ to calibrate functional CP band $C_{\alpha}$ and evaluate on $\mathcal{D}_{\text{eval-unseen}}$. By varying the user-specified $\alpha$, we can adjust the conservativeness of the failure detectors and obtain a trade-off between accuracy (bal-acc) and detection time (T-det). To be a good failure detector, it should detect failures both accurately (higher bal-acc) and promptly (lower T-det), and thus have the curve rise toward the top-left corner in the plots of Fig. 3. As we can see from Fig. 3, the proposed SAFE-MLP and SAFE-LSTM perform the best on OpenVLA+LIBERO and $\pi_{0}$+SimplerEnv benchmarks, and on par with the best baseline on the other two benchmarks. We also manually annotate the ground truth (GT) failure timesteps (when a human thinks that failure happens or intervention is needed) for failed rollouts, and plot them as blue vertical lines in Fig. 3. Comparing SAFE’s performance with the GT fail time, we can see that SAFE can detect failures with high accuracy in the early stages of rollouts and potentially before the failure happens. This early detection allows early intervention for policy failures before they get stuck in execution or cause harm to the real-world environment.
6.3 What are detected failure modes and are they aligned with human intuition?
In Fig. 4 and Fig. 5(b-d), we visualize rollouts with the failure scores detected by SAFE. Fig. 4 demonstrates common failure modes in simulation, including imprecise insertion, unstable actions and missed grasps. Two successful rollouts on the real robot are shown in Fig. 5(b-c), where failure scores stop increasing after task completion. For the failed rollouts, the failure flag is raised after the policy is frozen (Fig. 5d) or the object slips out of the gripper (Fig. 5e). This aligns well with human intuition. Please refer to Appendix for video illustrations.
7 Conclusion
In this paper, we introduce the multitask failure detection problem for generalist VLA policies, where failure detectors are trained only on seen tasks and evaluated on unseen tasks. We analyze VLA’s internal feature space and find that the internal features are separated for successful and failed rollouts. Based on this observation, we propose SAFE, a simple and efficient failure detection method by operating on the VLA’s internal features. SAFE is evaluated on multiple VLAs in both simulation and the real world, and compared with diverse baselines. Experiments show that SAFE achieves SOTA results in failure detection, and aligns with human intuition.
Limitations: Most recent VLAs have shown capabilities in handling diverse modalities, controlling diverse embodiments and learning latent actions from non-robotic action-less video data [66, 67]. This paper only considers multitask failure detection for manipulation tasks, and it’s not clear how well the failure detectors generalize across embodiments, sim2real or to the action-less videos. Besides, SAFE only uses features from the last layer and how to effectively aggregate information across multiple layers of a VLA remains future work.
Acknowledgments and Disclosure of Funding
The authors were partially supported by the Toyota Research Institute (TRI) and NSERC Discovery Grant. The authors thank Blerim Abdullai, Sebastian Aegidius, Jasper Gerigk, Ruthrash Hari, and Wei-Cheng Tseng for helpful discussions and feedback.
Appendix
Appendix A Experiment Details
A.1 Vision-Language-Action Models
We conduct experiments on 3 state-of-the-art large VLA models: OpenVLA [2], $\pi_{0}$ [4] and $\pi_{0}$-FAST [5]. Given the internal feature vectors $E\in\mathbb{R}^{n\times d^{\prime}}$ produced by a VLA model, where dimension $n$ corresponds to different token positions, diffusion steps, etc. and $d$ is the feature dimension, we aggregate $E$ into a single fixed-dimensional feature vector $\mathbf{e}\in\mathbb{R}^{d}$ before inputting to the proposed SAFE models. In this paper, we consider and ablate the following ways of feature aggregation:
  • First: take the first vector along the dimension $n$, $\mathbf{e}=E_{1}$;
  • Last: take the last vector along the dimension $n$, $\mathbf{e}=E_{n}$;
  • Mean: take average over the dimension $n$, $\mathbf{e}=\frac{1}{n}\sum_{i=1}^{n}E_{i}$;
  • First&Last: concatenate the first and the last vector, $\mathbf{e}=\text{concat}(E_{0},E_{n})\in\mathbb{R}^{2d^{\prime}}$.
Both OpenVLA and $\pi_{0}$-FAST first predict a sequence of discrete tokens and then convert them into continuous actions. We take feature vectors before being decoded into the output tokens from the last transformer block as $E\in\mathbb{R}^{n\times d^{\prime}}$, and therefore $n$ corresponds to the number of generated tokens. We ablate the aggregation method along this dimension and denote the aggregation method as $agg_{\text{token}}$. For $\pi_{0}$-FAST, we additionally ablate using the feature vectors before (“encoded”) and after (“pre-logits”) the final RMS normalization layer as $E$.
Differently, $\pi_{0}$ (and $\pi_{0}^{*}$) outputs action vectors by flow matching [68], and we take the feature vectors before being projected into the velocity field. Suppose $\pi_{0}$ predict an action chunk of horizon $H$ and performs $k$ flow matching steps, the internal features become $E\in\mathbb{R}^{H\times k\times d}$ and we perform the aggregation process along the $H$ dimension and the $k$ dimension separately to get the final embedding vector $\mathbf{e}\in\mathbb{R}^{d^{\prime}}$. The aggregation methods are denoted as $agg_{\text{hori}}$ and $agg_{\text{diff}}$ along these two dimensions, respectively.
For all VLA models, we ablate different methods of aggregating the hidden features $E$ into a single feature vector $\mathbf{e}$ and select the best method according to $\mathcal{D}_{\text{eval-seen}}$ performance. The detailed ablation results are shown in Section A.7.
OpenVLA and $\pi_{0}^{*}$ use MIT License; $\pi_{0}$ and $\pi_{0}$-FAST use Apache-2.0 license.
A.2 SAFE Failure Detector
SAFE-LSTM uses an LSTM model with 1 layer and a hidden dimension of 256, and an additional linear layer is used to project the hidden states of LSTM into a single scalar $s_{t}$. SAFE-MLP uses a multi-layer perceptron with 2 layers and a hidden dimension of 256. Since successes and failures from the generated rollouts are imbalanced, the losses on positive (failed) and negative (successful) rollouts are weighted by their inverse class frequency. We also apply an L2 regularization loss on the model weights to reduce overfitting, and this loss is weighted by $\lambda_{\text{reg}}$ and optimized together with the failure score learning loss $L_{\text{LSTM}}$ or $L_{\text{MLP}}$. $\lambda_{\text{reg}}$ are determined by grid search.
A.3 Failure Detection Baselines
For the cluster entropy baseline, we use agglomerative clustering with the ward linkage criterion [69]. The distance threshold is denoted as $\delta$ and decided by grid search.
For the STAC baseline [17], we use the Maximum Mean Discrepancy (MMD) distance measure [70] with radial basis function kernels, which was reported to have the best performance by [17]. The bandwidth of the RBF kernel is $1$.
For all baselines except for RND [39] and LogpZO [8], we ablate one version that only considers the failure score computed from the current timestep (“cumsum=False”) and another that uses the cumulative sum (cumsum) of the failure scores over time (“cumsum=True”).
For RND and LogpZO, we use the original implementation provided by the authors  [^3] https://github.com/CXU-TRI/FAIL-Detect and do not accumulate scores. In [8], RND and LogpZO are trained to model the distribution of (encoded) observations $\mathbf{o}_{t}$ and predicted actions $\mathbf{A}_{t}$. In this work, we adapt them to model the distribution of VLA’s internal embeddings $\mathbf{e}_{t}$.
Note that as $\pi_{0}$ (and $\pi_{0}^{*}$) does not output discrete tokens, token uncertainty-based baselines do not apply. And for OpenVLA, $H=H^{\prime}=1$ and thus the STAC [17] and STAC-Single do not apply.
A.4 Conformal Prediction
We follow [22, 8] for CP band construction. Please refer to Section. B in the Appendix of [8] for a detailed formulation. Specifically, in our experiments, we use the adaptive modulation function (Equation 2 in the Appendix of [8]), which models the non-extreme behaviors of the functional data.
A.5 Benchmark Details
LIBERO [20]: We adopt the LIBERO-10 task suite, which contains the most diverse set of objects, environments, and instructions among the 4 LIBERO task suites, and therefore LIBERO-10 is considered the most challenging task suite. LIBERO-10 contains 10 tasks with 50 rollouts in each task. We use the initial conditions for all rollouts as specified and provided by the author [^4] https://github.com/Lifelong-Robot-Learning/LIBERO/tree/master/libero/libero/init_files. To test VLA models on LIBERO, we adopt the trained model weights provided by the respective authors and do not further finetune them. On LIBERO-10, OpenVLA achieves a success rate of 53.7%, $\pi_{0}$-FAST achieves 60.2%, and $\pi_{0}$ achieves 85.2%. For evaluation, 3 out of 10 tasks are unseen, and within seen tasks, 60% of rollouts are used for $\mathcal{D}_{\text{train}}$ and the remaining 40% for $\mathcal{D}_{\text{eval-seen}}$.
Note that the LIBERO simulator stops the rollout execution when the robot finishes the task (considered a success) or a maximum rollout length is reached (considered a failure). Therefore, in the generated rollouts, failed ones always have the maximum length, but successful ones are shorter. This could result in an unfair advantage for some of the compared failure detectors (if a failure detector simply learns to count the time elapsed, i.e., $s_{t}=t$, it will achieve perfect failure detection since failed rollouts have a fixed and longer duration). To ensure a fair comparison, for evaluation in Table 1, we compute the minimum rollout length for each task and use that as $T$ for that task. The failure detection performance (in ROC-AUC) is then determined based on $s_{T}$, where $T$ is the same for all successful and failed rollouts within each task.
LIBERO benchmark uses the MIT license.
SimplerEnv [61]: SimplerEnv carefully identifies and reduces the domain gap between the simulation and the real-world demonstration data, and provides simulated environments that highly resemble the demonstration data from RT-series [1, 7, 32] (with the Google Robot embodiment) and BridgeData V2 [34] (with the WidowX embodiment). They show that models pretrained on real-world datasets can also accomplish similar tasks in SimplerEnv without finetuning, and their performance in simulation matches that in the real world.
On this benchmark, we adopt the pretrained model checkpoints of $\pi_{0}^{*}$ [62]. Note that $\pi_{0}^{*}$ model checkpoints are trained separately on the Google Robot embodiment and the WidowX embodiment, which results in two model checkpoints that have different internal feature spaces. Therefore, all failure detectors are trained and evaluated on each embodiment separately as well. All reported evaluation metrics are computed separately for each embodiment and then averaged. In Table 2, we list the tasks used for failure detection on SimplerEnv. We generate 100 rollouts for each task with random initial configurations, and the success rates of $\pi_{0}^{*}$ on each task are also listed in Table 2. A rollout stops after the maximum number of allowed timesteps have passed, regardless of task success or failure. Within each embodiment, 1 out of 4 tasks is unseen, and within the seen tasks, 66% of the rollouts are in $\mathcal{D}_{\text{train}}$ and the remaining 33% in $\mathcal{D}_{\text{eval-seen}}$.
SimplerEnv benchmark uses the MIT license.
Table 2: List of tasks used in SimplerEnv benchmark.
Embodiment Task ID Environment Name $\pi_{0}^{*}$ Success Rate (%) Google Robot 1 google_robot_move_near_v0 77 Google Robot 2 google_robot_open_drawer 50 Google Robot 3 google_robot_close_drawer 80 Google Robot 4 google_robot_place_apple_in_closed_top_drawer 40 WidowX 1 widowx_carrot_on_plate 44 WidowX 2 widowx_put_eggplant_in_basket 88 WidowX 3 widowx_spoon_on_towel 79 WidowX 4 widowx_stack_cube 43
Table 3: List of tasks used in the real-world experiments.
Task Instruction Rollout Length $T$
1 close the door 300
2 close the drawer 200
3 pick up the ball and place it in the bowl 400
4 pick up the knife and put it on the plate 350
5 pick up the lid and place it on the pot 400
6 pick up the lid from the pot and place it on the table 400
7 pick up the marker and place it in the cup 400
8 place the green block on the yellow block 350
9 place the pink cup to the right of the blue cup 300
10 press the button 200
11 put both the carrot and the ball in the bowl 500
12 put the cup to the upright position 500
13 unfold the cloth 500
Real-world experiments with Franka robot: In Table 3, we list the tasks used in the real-world experiments. For each task, we set a number of timesteps $T$ allowed for one rollout, and all rollouts of the same task are terminated after the same $T$ timesteps regardless of task success or failure. In Fig. 6, we further visualize some example successful and failed rollouts from the real-world experiments.
Figure 6: Example successful and failed rollouts from real-world experiments.
A.6 Training Details
We use Adam optimizer [71] with $\beta_{1}=0.9$, $\beta_{1}=0.999$, $\epsilon=10^{-8}$, and a learning rate (lr) determined by grid search. The SAFE models are trained for 1000 epochs with batch size 512. Note that each rollout is considered as one data point and thus batch size of 512 translates to training on (at most) 512 rollouts in each iteration. All training and evaluation are done on a single NVIDIA A100 40GB GPU. Since SAFE uses small networks (MLP or LSTM with 1 or 2 layers), the typical training time for one model is less than one minute.
A.7 Hyperparameter Tuning
To determine the hyperparameters for the proposed SAFE and baselines, we perform a grid search over them and select the ones with the highest failure detection performance (ROC-AUC) on the $\mathcal{D}_{\text{eval-seen}}$ split. In Table 5, Table 6, and Table 7, we report the hyperparameters we have searched over and the values with the best performance. Note that for the real-world experiments, we fix the $\mathbf{e}_{t}$ to be the “pre-logits” with “Mean” aggregation.
Appendix B Additional Results and Discussions
B.1 Feature Visualization and Analysis
We perform the feature analysis similar to Section 4.1 and Fig. 1 on other benchmarks and show the plots in Fig. 7, Note that in this feature analysis process, the t-SNE algorithm was performed on the VLA’s embeddings without any learning. Therefore, the feature dimension reduction process is unsupervised and does not know about task successes or failures.
Comparing the plots in Fig. 1 and Fig. 7, we can see that the embedding spaces from VLAs are different from each other, which corresponds to the different failure patterns presented by the VLAs. For $\pi_{0}$-FAST (Fig. 1) and $\pi_{0}$ on LIBERO (Fig. 7a and b), when task execution fails, the embeddings fall into the same region (“failure zone”). This corresponds to the major failure mode of $\pi_{0}$-FAST trained on the LIBERO dataset, where the predicted actions $\mathbf{A}_{t}$ become unstable and the robot arms move to weird configurations and out of the observation frame. For OpenVLA on LIBERO (Fig. 7c and d). we observe that for most failed rollouts, the robot freezes at or shakes around certain configurations during the middle of task execution. Such failed rollouts result in features very close to each other, which corresponds to small blobs of red dots in Fig. 7c.
Despite the different appearances of the embedding spaces from the above benchmarks, their successful and failed rollouts are separable in the feature space. This is aligned with the high performance of the proposed SAFE and the embedding-based baseline methods. Moreover, although the embeddings of the failed rollouts from OpenVLA are spread over the space and do not form a unified “failure zone”, SAFE is still able to learn to separate task failures from successes (possibly by extracting the correlations that are not visualized by t-SNE) and generalize well to unseen tasks, as reported in Table 1.
However, the visualized embeddings of $\pi_{0}$-FAST on the real Franka robot (Fig. 7e and f) are different, where embeddings from successful and failed rollouts are not easily separable through the t-SNE visualization. We hypothesize that because the tasks we used for real-world experiments are more diverse, their failures do not have a unified semantic meaning, and thus the embeddings are not clearly separated in the visualization. This explains the limited performance of all failure detection methods as reported by Fig. 5, where ROC-AUC is at most 64 on $\mathcal{D}_{\text{eval-unseen}}$. Nevertheless, SAFE-MLP still outperforms all baselines on both seen and unseen splits in this evaluation.
(a) $\pi_{0}$+LIBERO, colored by task success
(b) $\pi_{0}$+LIBERO, colored by task ID
(c) OpenVLA+LIBERO, colored by task success
(d) OpenVLA+LIBERO, colored by task ID
(e) $\pi_{0}$-FAST+Franka, colored by task success
(f) $\pi_{0}$-FAST+Franka, colored by task ID
Figure 7: t-SNE plots of VLA’s internal features, from different evaluation benchmarks.
B.2 Conformal Prediction Results
We use functional CP [8, 22] to determine the time-varying thresholds $\delta_{t}$ for failure detection. By varying the significance level $\alpha$ used in functional CP, we can adjust the conservativeness of failure detection and get different performance. In Fig. 8, we plot the change of TNR (True Negative Rate), TPR (True Positive Rate) and Bal-acc (Balanced Accuracy, $\frac{\text{TNR}+\text{TPR}}{2}$) w.r.t. $\alpha$.
Note that we calibrate the CP bands on successful rollouts (negative data points), and thus if the assumptions used in CP ($s_{t}$ are sampled i.i.d.) hold, the TNR rate is lower bounded by and close to $1-\alpha$ (the gray dashed line in the TNR plots in Fig. 8[22]. However, as the multitask failure detection problem requires detecting failures on tasks that are not in the training or the calibration sets, we need to calibrate CP bands on $\mathcal{D}_{\text{eval-seen}}$ and then evaluate them on $\mathcal{D}_{\text{eval-unseen}}$. Therefore, the i.i.d assumption may not hold, and TNR may deviate from the gray dashed line.
From Fig. 8, we can see that on OpenVLA+LIBERO and $\pi_{)}^{*}$+SimplerEnv benchmarks, the TNR curves obtained by SAFE are close to the gray dashed line $1-\alpha$, while those on the other 3 benchmarks are lower than $1-\alpha$. A similar phenomenon is also observed for the baseline methods: none of the TNR curves obtained from the baselines consistently conform to the $1-\alpha$ curve across all benchmarks. We attribute this to the challenging nature of the multitask failure detection problem, where the failure scores for calibration and evaluation may not come from the same distribution. Nevertheless, we still adopt the functional CP as a principled method to determine the time-varying failure detection threshold $\delta_{t}$. Moreover, from Fig. 8, we can see that SAFE can achieve higher TPR and result in fewer false negatives compared to the baselines. This is crucial for safety-critical environments, where a missing failure (false negative) can be much more catastrophic than a false alarm (false positive).
Figure 8: Additional failure detection results using $\delta_{t}$ obtained by functional CP. These plots show TNR (left column), TPR (middle column), and Bal-acc (right column) w.r.t. the significance level $\alpha$, for each evaluation benchmark. These plots are obtained with random seed$=0$.
B.3 Failure Detection Time
As mentioned in Section 6.2, we manually label when the failure happens for the failed rollouts. The labeling process is based on video recordings after all rollouts are collected and no interventions were done during the task execution. While the exact times of failure are clear for some failure modes (e.g. dangerous actions, breaking objects), they can be ambiguous and hard to annotate for other failure modes. For example, a policy may freeze in the middle of task execution, and after that either recovering from it or getting stuck indefinitely can be possible. In another case, a policy may repeatedly try grasping the object but keep missing the grasp until timeout, and it’s hard to determine a single point of failure. To handle such cases, we instruct the human annotators to pick the time where they think intervention is needed and they should take over control to prevent an execution failure. In practice, for the above ambiguous failure modes, we annotate the failures after the policy gets stuck by a few seconds or re-tries the grasping action a few times. For some rollouts that look very plausible but do not succeed due to the time limit, the failure time is annotated as the end of the rollout. Note that we annotate only the failed rollouts and not the successful ones, even though they may also show subtle signs of failure in the middle.
In Fig. 9, we compare the times of failure detected by the proposed SAFE-MLP model and a human annotator. From Fig. 9, we can see that for both $\pi_{0}$ and $\pi_{0}$-FAST models, SAFE-MLP can detect failures before they happen (as identified by a human). When used for $\pi_{0}$-FAST deployed on LIBERO, SAFE-MLP can forecast failures well in advance and even predict 40% of the failures after the first timestep.
Furthermore, from Fig. 9a and Fig. 9c, we can see that the blue curves jump up on the right edge of the plots. This means that the human annotator does not think these rollouts are failures until the very last moment, where the VLA model is probably on the right track and fails only due to timeout. We think such failures are also hard for failure detectors to detect, and it explains the low performance of all failure detectors on these benchmarks.
(a) $\pi_{0}$ on LIBERO benchmark.
(b) $\pi_{0}$-FAST on LIBERO benchmark.
(c) $\pi_{0}$-FAST on the real Franka robot.
Figure 9: Comparison between detected and ground truth (GT) failure w.r.t time. On the left column, we plot the cumulative number of true failures (true positives) detected by SAFE-MLP (red) and a human annotator (blue), w.r.t. elapsed time in each rollout. The right column shows the time of failures detected by SAFE-MLP (y-axis) and a human annotator (x-axis) for each rollout, where failures missed by the detector (false negatives) are plotted in blue crosses. Experiments are done with seed 0 and functional CP with significance level $\alpha=0.15$.
B.4 Result Variance
In Table 4, we report the standard deviation for all results in Table 1 and Fig. 5 left. Note that for the repeated runs, not only are they using different random seeds, also the tasks are split differently into the seen and the unseen subsets. Since different tasks have different difficulties for failure detection, it is normal to see large standard deviations in Table 4. From Table 4, we can see that the proposed SAFE methods achieve high averaged performance with relatively low standard deviations compared to the baselines, across all evaluation benchmarks. This signifies the strong and also stable performance of SAFE.
Table 4: Mean and standard deviation of failure detection ROC-AUC on all benchmarks. This table complements the results from Table 1 and Fig. 5 left.
VLA Model OpenVLA $\pi_{0}$-FAST $\pi_{0}$ $\pi_{0}^{*}$ $\pi_{0}$-FAST Benchmark LIBERO LIBERO LIBERO SimplerEnv Real Franka Eval Task Split Seen Unseen Seen Unseen Seen Unseen Seen Unseen Seen Unseen Max prob. 50.25$\pm$2.51 53.83$\pm$6.32 61.32$\pm$9.57 69.44$\pm$13.61 - - - - 53.74$\pm$3.46 48.59$\pm$3.00 Avg prob. 44.05$\pm$1.26 51.58$\pm$1.82 52.46$\pm$3.44 58.04$\pm$5.64 - - - - 51.60$\pm$3.12 47.30$\pm$4.32 Max entropy 52.94$\pm$4.36 53.09$\pm$7.68 46.69$\pm$13.33 62.96$\pm$19.62 - - - - 59.23$\pm$3.06 53.50$\pm$3.15 Avg entropy 45.27$\pm$1.78 50.03$\pm$3.18 50.93$\pm$1.22 58.63$\pm$3.47 - - - - 50.67$\pm$3.96 46.08$\pm$4.79 Mahalanobis dist. 62.03$\pm$5.11 58.85$\pm$4.16 93.56$\pm$2.32 83.79$\pm$7.18 77.12$\pm$8.57 74.31$\pm$12.64 88.42$\pm$2.82 52.84$\pm$31.97 75.54$\pm$4.07 53.93$\pm$5.06 Euclidean dist. $k$-NN 66.00$\pm$2.33 55.23$\pm$10.05 92.04$\pm$2.39 84.12$\pm$6.47 75.64$\pm$6.20 70.73$\pm$16.69 89.73$\pm$3.08 68.41$\pm$9.22 80.35$\pm$5.36 60.27$\pm$4.79 Cosine dist. $k$-NN 67.09$\pm$2.74 69.45$\pm$6.14 92.09$\pm$1.70 84.64$\pm$4.90 75.76$\pm$6.16 70.31$\pm$16.84 90.19$\pm$4.05 71.32$\pm$12.02 80.23$\pm$5.12 59.51$\pm$5.76 PCA-KMeans [9] 57.18$\pm$2.04 55.10$\pm$1.16 68.46$\pm$4.92 57.12$\pm$10.44 64.92$\pm$8.90 60.35$\pm$19.93 66.88$\pm$5.10 61.19$\pm$14.76 51.91$\pm$4.20 49.86$\pm$6.19 RND [39] 52.57$\pm$4.56 46.88$\pm$4.92 88.67$\pm$3.05 81.57$\pm$8.67 71.92$\pm$7.02 69.44$\pm$19.39 85.07$\pm$4.04 65.89$\pm$6.52 62.00$\pm$5.44 45.83$\pm$5.10 LogpZO [8] 61.57$\pm$3.62 52.91$\pm$5.79 91.52$\pm$2.39 83.07$\pm$7.17 76.80$\pm$9.12 73.23$\pm$11.64 88.79$\pm$4.92 74.66$\pm$14.96 64.43$\pm$7.82 52.24$\pm$3.68 Action total var. 62.76$\pm$1.66 65.43$\pm$2.50 76.95$\pm$7.22 74.50$\pm$12.19 77.20$\pm$5.65 75.18$\pm$5.08 68.41$\pm$10.81 67.94$\pm$15.97 - - Trans. total var. 55.33$\pm$2.06 58.99$\pm$5.13 78.21$\pm$4.09 80.03$\pm$9.11 49.38$\pm$9.95 54.71$\pm$7.57 63.27$\pm$7.17 55.90$\pm$19.19 - - Rot. total var. 47.85$\pm$2.88 55.30$\pm$4.38 80.87$\pm$5.85 77.29$\pm$8.71 52.94$\pm$7.56 61.06$\pm$10.60 58.07$\pm$10.41 62.10$\pm$9.39 - - Gripper total var. 61.84$\pm$2.67 64.48$\pm$3.05 76.82$\pm$7.10 74.42$\pm$12.13 77.19$\pm$5.66 75.19$\pm$5.08 69.16$\pm$9.50 69.29$\pm$14.77 - - Cluster entropy 50.16$\pm$2.36 51.44$\pm$1.01 80.22$\pm$7.37 80.53$\pm$8.65 76.19$\pm$4.31 72.12$\pm$1.04 68.25$\pm$9.03 73.66$\pm$16.03 - - STAC [17] - - 83.07$\pm$4.61 85.31$\pm$6.71 46.55$\pm$8.90 47.91$\pm$20.94 60.74$\pm$13.89 62.21$\pm$16.72 - - STAC-Single - - 85.46$\pm$6.55 81.16$\pm$8.63 68.46$\pm$5.10 69.39$\pm$8.22 68.71$\pm$7.06 70.40$\pm$8.76 45.24$\pm$3.68 38.01$\pm$9.81 SAFE-LSTM 70.24$\pm$1.49 72.47$\pm$5.55 92.98$\pm$2.62 84.48$\pm$7.29 76.98$\pm$5.34 71.09$\pm$6.94 88.85$\pm$6.30 80.11$\pm$10.49 77.27$\pm$5.82 58.70$\pm$4.37 SAFE-MLP 72.68$\pm$2.38 73.47$\pm$5.39 90.06$\pm$2.82 80.44$\pm$5.72 73.50$\pm$7.43 73.27$\pm$11.85 89.50$\pm$4.49 84.82$\pm$8.12 86.76$\pm$2.64 64.16$\pm$5.88
Table 5: Grid-searched and best-performing hyperparameters (in bold text) for OpenVLA+LIBERO (left) and $\pi_{0}$-FAST+LIBERO (right).
Method HParams Values Max prob. cumsum True False Avg prob. cumsum True False Max entropy cumsum True False Avg entropy cumsum True False Mahalanobis dist. $agg_{\text{token}}$ First Last Mean cumsum True False Euclidean dist. $k$-NN $agg_{\text{token}}$ First Last Mean cumsum True False $k$ 1 5 10 Cosine dist. $k$-NN $agg_{\text{token}}$ First Last Mean cumsum True False $k$ 1 5 10 PCA-KMeans $agg_{\text{token}}$ First Last Mean cumsum True False clusters 16 32 64 dim 32 64 128 RND $agg_{\text{token}}$ First Last Mean LogpZO $agg_{\text{token}}$ First Last Mean Action total var. cumsum True False Trans. total var. cumsum True False Rot. total var. cumsum True False Gripper total var. cumsum True False Cluster entropy cumsum True False $\delta$ 0.01 0.05 SAFE-LSTM $agg_{\text{token}}$ First Last Mean lr 1e-4 3e-4 1e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1 1 SAFE-MLP $agg_{\text{token}}$ First Last Mean lr 1e-4 3e-4 1e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1 1
Method HParams Values Max prob. cumsum True False Avg prob. cumsum True False Max entropy cumsum True False Avg entropy cumsum True False Mahalanobis dist. $agg_{\text{token}}$ First Last Mean Feat Encoded Pre-logits cumsum True False Euclidean dist. $k$-NN $agg_{\text{token}}$ First Last Mean Feat Encoded Pre-logits cumsum True False $k$ 1 5 10 Cosine dist. $k$-NN $agg_{\text{token}}$ First Last Mean Feat Encoded Pre-logits cumsum True False $k$ 1 5 10 PCA-KMeans $agg_{\text{token}}$ First Last Mean Feat Encoded Pre-logits cumsum True False clusters 16 32 64 dim 32 64 128 RND $agg_{\text{token}}$ First Last Mean Feat Encoded Pre-logits LogpZO $agg_{\text{token}}$ First Last Mean Feat Encoded Pre-logits Action total var. cumsum True False Trans. total var. cumsum True False Rot. total var. cumsum True False Gripper total var. cumsum True False Cluster entropy cumsum True False $\delta$ 0.01 0.05 0.1 0.2 0.5 1 2 5 STAC cumsum True False STAC-Single cumsum True False SAFE-LSTM $agg_{\text{token}}$ First Last Mean Feat Encoded Pre-logits lr 3e-5 1e-4 3e-4 1e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1 SAFE-MLP $agg_{\text{token}}$ First Last Mean Feat Encoded Pre-logits lr 3e-5 1e-4 3e-4 1e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1
Table 6: Grid-searched and best-performing hyperparameters (in bold text) for $\pi_{0}$+LIBERO (left) and $\pi_{0}^{*}$+SimplerEnv (right).
Method HParams Values Mahalanobis dist. $agg_{\text{hori}}$ First Last First&Last $agg_{\text{diff}}$ First Last First&Last cumsum True False Euclidean dist. $k$-NN $agg_{\text{hori}}$ First Last First&Last $agg_{\text{diff}}$ First Last First&Last cumsum True False $k$ 1 5 10 Cosine dist. $k$-NN $agg_{\text{hori}}$ First Last First&Last $agg_{\text{diff}}$ First Last First&Last cumsum True False $k$ 1 5 10 PCA-KMeans $agg_{\text{hori}}$ First Last First&Last $agg_{\text{diff}}$ First Last First&Last cumsum True False clusters 16 32 64 dim 32 64 128 RND $agg_{\text{hori}}$ First Last First&Last $agg_{\text{diff}}$ First Last First&Last LogpZO $agg_{\text{hori}}$ First Last First&Last $agg_{\text{diff}}$ First Last First&Last Action total var. cumsum True False Trans. total var. cumsum True False Rot. total var. cumsum True False Gripper total var. cumsum True False Cluster entropy cumsum True False $\delta$ 0.01 0.05 0.1 0.2 0.5 1 2 5 STAC cumsum True False STAC-Single cumsum True False SAFE-LSTM $agg_{\text{hori}}$ First Last First&Last $agg_{\text{diff}}$ First Last First&Last lr 1e-5 3e-5 1e-4 3e-4 1e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1 SAFE-MLP $agg_{\text{hori}}$ First Last First&Last $agg_{\text{diff}}$ First Last First&Last lr 1e-5 3e-5 1e-4 3e-4 1e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1
Method HParams Values Mahalanobis dist. $agg_{\text{hori}}$ First Last Mean First&Last $agg_{\text{diff}}$ First Last Mean First&Last cumsum True False Euclidean dist. $k$-NN $agg_{\text{hori}}$ First Last Mean First&Last $agg_{\text{diff}}$ First Last Mean First&Last cumsum True False $k$ 1 5 10 Cosine dist. $k$-NN $agg_{\text{hori}}$ First Last Mean First&Last $agg_{\text{diff}}$ First Last Mean First&Last cumsum True False $k$ 1 5 10 PCA-KMeans $agg_{\text{hori}}$ First Last Mean First&Last $agg_{\text{diff}}$ First Last Mean First&Last cumsum True False clusters 16 32 64 dim 32 64 128 RND $agg_{\text{hori}}$ First Last Mean First&Last $agg_{\text{diff}}$ First Last Mean First&Last LogpZO $agg_{\text{hori}}$ First Last Mean First&Last $agg_{\text{diff}}$ First Last Mean First&Last Action total var. cumsum True False Trans. total var. cumsum True False Rot. total var. cumsum True False Gripper total var. cumsum True False Cluster entropy cumsum True False $\delta$ 0.01 0.05 0.1 0.2 0.5 1 2 5 STAC cumsum True False STAC-Single cumsum True False SAFE-LSTM $agg_{\text{hori}}$ First Last Mean First&Last $agg_{\text{diff}}$ First Last Mean First&Last lr 1e-4 3e-4 1e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1 1 SAFE-MLP $agg_{\text{hori}}$ First Last Mean First&Last $agg_{\text{diff}}$ First Last Mean First&Last lr 1e-4 3e-4 1e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1 1
Table 7: Grid-searched and best-performing hyperparameters (in bold text) for $\pi_{0}$-FAST on real-world rollouts.
Method HParams Values Max prob. cumsum True False Avg prob. cumsum True False Max entropy cumsum True False Avg entropy cumsum True False Mahalanobis dist. cumsum True False Euclidean dist. $k$-NN cumsum True False $k$ 1 5 10 Cosine dist. $k$-NN cumsum True False $k$ 1 5 10 PCA-KMeans cumsum True False clusters 16 32 64 dim 32 64 128 STAC-Single cumsum True False SAFE-LSTM lr 1e-4 3e-4 1e-3 3e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1 SAFE-MLP lr 1e-4 3e-4 1e-3 3e-3 $\lambda_{reg}$ 1e-3 1e-2 1e-1